OpenStudy (anonymous):

y^2+6y+8x+25=0 find the vertex focus and directrix of the parabola

OpenStudy (anonymous):

y^2+6y+8x+25=0 (y+3)^2=-8(x+2) vertex=(-2,-3) focus=(-2,-1) and directrix=y+3=-2;

OpenStudy (anonymous):

the focus (-4,-3) and the directrix is x=0 but how did u even go about figuring the prblem out